Calculate the heat of reaction: 2 NO(g)+O2(g)-->2NO2(g)


1
Expert's answer
2021-05-18T06:20:02-0400

remaining information

1/2N2 (g) + O2 (g) → NO2 (g), ΔH°A = 33.2 kJ .....(1)

1/2N2 (g) + 1/2O2 (g) → NO (g), ΔH°B = 90.2 kJ ....(2)


multiplying equation (1) by 2 and in equation 2 , first multiply by 2 and reverse it and then add both the equations

we get required equation


 2 NO(g)+O2(g)-->2NO2(g)

final ="(32.2\times2-90.2\times)kJ=114kJ"